Tom wants to build a fence around his rectangular swimming pool. The dimensions of the pool are 18 feet by 22feet. He wants the border around the pool to be of uniform width. If he has 60 yards of fencing, find the width of the border.
Answer by nerdybill(7384)

Tom wants to build a fence around his rectangular swimming pool. The dimensions of the pool are 18 feet by 22feet. He wants the border around the pool to be of uniform width. If he has 60 yards of fencing, find the width of the border.
.
Convert 60 yards to feet:
60 * 3 = 180 feet
.
Let x = width of border
then
18+2x = width of fence
22+2x = length of fence
.
perimeter = 2(width+length)
180 = 2(18+2x + 22+2x)
180 = 2(40+4x)
90 = 40+4x
50 = 4x
12.5 feet = x